It takes 3 hrs for a steamer to travel 27 kilometers upstream. The same steamer can travel 30 kilometers downstream in 2 hrs. Determine the speed of the steamer.

Answer by mananth(16949)

boat speed= x mph
current speed= y mph
with stream speed= (x+y)
against stream speed = (x-y)
d= 30
30 / (x+y) = 2
divide by 2
15 / (x+y) = 1
(x+y) = 15 ............1
d= = 27
27/(x-y) = 3
divide by 3
27 /(x-y) = 3
x-y = 9 .............2
add up (1) & (2)
2 x = 24
/ 2
x= 12 mph speed of boat in still water
plug value of x in (1)
we get y= 3 mph speed of current
m.ananth@hotmail.ca